# Continual Learning: Forget-free Winning Subnetworks for Video Representations

### v3

@@ 3.3.2 Soft-Subnetworks (SoftNet) with FSO
3.4 Optimization for TIL, VIL, and FSCIL
3.4.1 Winning SubNetworks (WSN) for TIL
@@
Algorithm 1 Winning SubNetworks (WSN) for TIL and VIL.
@@ 3.4.3 Soft-SubNetworks (SoftNet) for FSCIL
17:       $\bm{\theta}_{\ast}\leftarrow\bm{\theta}_{\ast}-\beta\left(\frac{\partial\mathcal{L}}{\partial\bm{\theta}_{\ast}}\odot\bm{m}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}minor}}\right)$
@@ 4 Experiments
We validate our method on several benchmark datasets against relevant continual learning baselines on Task-Incremental Learning (TIL, Section 3.4.1), Video Incremental Learning (VIL, Section 3.4.2), and Few-shot Class Incremental Learning (FSCIL, Section 3.4.3).
@@ 4.1 Task-incremental Learning (TIL)
Experimental settings. As we directly implement our method from the official code of [68], we provide the values for HAT and GPM reported in [68]. For Omniglot Rotation and Split CIFAR-100 Superclass, we deploy the proposed architecture in multi-head settings with hyperparameters as reported in [95]. All our experiments run on a single-GPU setup of NVIDIA V100.
Performance metrics. We evaluate all methods based on the following four metrics:
@@
    Accuracy (ACC) measures the average of the final classification accuracy on all tasks: $\mathrm{ACC}=\frac{1}{T}\sum_{i=1}^{T}A_{T,i}$, where $A_{T,i}$ is the test accuracy for task $i$ after training on task $T$.
@@
    Backward Transfer (BWT) measures the forgetting during continual learning. Negative BWT means that learning new tasks causes the forgetting of past tasks: $\mathrm{BWT}=\frac{1}{T-1}\sum_{i=1}^{T-1}A_{T,i}-A_{i,i}$.
@@ 4.2 Few-shot Class Incremental Learning (FSCIL)
Table I: (TIL), Performance comparisons of the proposed method and other baselines - PackNet [54] and SupSup [90] - on various benchmark datasets. We report the mean and standard deviation of the average accuracy (ACC) and average backward transfer (BWT) across $5$ independent runs with five seeds under the same experimental setup [17]. normal-$\dagger$    denotes results reported from [17].
@@
La-MaML [22] 71.37 ($\pm$ ${}^{\dagger}$5.39 ($\pm$ ${}^{\dagger}$ 54.44 ($\pm$ ${}^{\dagger}$6.65 ($\pm$ ${}^{\dagger}$ 66.90 ($\pm$ ${}^{\dagger}$9.13 ($\pm$ ${}^{\dagger}$
GPM [68] 73.18 ($\pm$ ${}^{\dagger}$1.17 ($\pm$ ${}^{\dagger}$ 57.33 ($\pm$ ${}^{\dagger}$0.37 ($\pm$ ${}^{\dagger}$ 67.39 ($\pm$ ${}^{\dagger}$  1.45 ($\pm$ ${}^{\dagger}$
FS-DGPM [17] 74.33 ($\pm$ ${}^{\dagger}$2.71 ($\pm$ ${}^{\dagger}$ 58.81 ($\pm$ ${}^{\dagger}$2.97 ($\pm$ ${}^{\dagger}$ 70.41 ($\pm$ ${}^{\dagger}$2.11 ($\pm$ ${}^{\dagger}$
@@
Multitask   79.75 ($\pm$ ${}^{\dagger}$   61.00 ($\pm$ ${}^{\dagger}$   77.10 ($\pm$ ${}^{\dagger}$
@@ 4.3 Video incremental Learning (VIL)
$\begin{split}\bm{\Gamma}(s,t)=&[\leavevmode\nobreak\ \sin(b^{0}\pi s),\cos(b^{0}\pi s),\cdots,\sin(b^{l-1}\pi s),\cos(b^{l-1}\pi s),\\ &\leavevmode\nobreak\ \leavevmode\nobreak\ \sin(b^{0}\pi t),\cos(b^{0}\pi t),\cdots,\sin(b^{l-1}\pi t),\cos(b^{l-1}\pi t)\leavevmode\nobreak\ ],\end{split}$ (11)
where the hyperparameters are set to $b=1.25$ and $l=80$ such that $\bm{\Gamma}(s,t)\in\mathbb{R}^{1\times 160}$. As differences from the previous NeRV model, the first layer of the MLP has its input size expanded from 80 to 160 to incorporate both frame and video indices, and distinct head layers after the NeRV block are utilized for each video. For the loss objective in Equation 8, $\alpha$ is set to $0.7$. We evaluate the video quality, average video session quality, and backward transfer with two metrics: PSNR.
@@
VIL’s performance metrics of PSNR We evaluate all methods based on the following continual learning metrics:
@@
    Average PSNR measures the average of the final performances on all video sessions: $\mathrm{PSNR}=\frac{1}{N}\sum_{s=1}^{N}A_{N,s}$, where $A_{N,s}$ is the test PSNR for session $s$ after training on the final video session $S$.
@@ 5.1 Comparisons with baselines in TIL
We use a multi-head setting to evaluate our WSN algorithm under the more challenging visual classification benchmarks. The WSN+FSO’s performances are compared with others in terms of two measurements on three major benchmark datasets as shown in Table I. Our WSN+FSO outperformed all state-of-the-art, achieving the best average accuracy of 79.00%, 61.70%, and 72.04%. WSN+FSO is also a forget-free model (BWT = ZERO), aligned with architecture-based models such as PackNet, SupSup, and WSN in these experiments. In addition, to show the effectiveness of the large single-scale task performance, we prepare WSN+FSO trained on the ImageNet-1K dataset, as shown in Table V. WSN+FSO outperformed all baselines.
@@ 6 Results On Few-shot CIL (FSCIL)
Table II: (FSCIL), Classification accuracy of ResNet18 on CIFAR-100 for 5-way 5-shot incremental learning with the same class split as in TOPIC [14]. ${}^{\ast}$ denotes the results reported from [72].
@@
Table III: (FSCIL), Classification accuracy of ResNet18 on miniImageNet for 5-way 5-shot incremental learning with the same class split as in TOPIC [14]. ${}^{\ast}$ denotes results reported from [72].
@@
Table IV: (FSCIL), Classification accuracy of ResNet18 on CUB-200-2011 for 10-way 5-shot incremental learning (TOPIC class split [82]). ${}^{\ast}$ denotes results reported from [72].
@@
Comparisons with SOTA. We compare SoftNet+FSO with the following state-of-art-methods on TOPIC class split [82] of three benchmark datasets - CIFAR100 (Table II), miniImageNet (Table III), and CUB-200-2011 (Table IV). Leveraged by regularized backbone ResNet, SoftNet+FSO outperformed all existing architecture baselines FSLL [55], FACT [100], and WaRP [37] on CIFAR100, miniImageNet. On CUB-200-201, the performances of SoftNet+FSO are comparable with those of ALICE and LIMIT, considering that ALICE used class/data augmentations and LIMIT added an extra multi-head attention layer. Note NC-FSCIL [94] could not be comparable with SoftNet+FSO since it focuses mainly on replaying prototype-based classifiers rather than backbone representations to obtain balanced categorical prototypes. Lastly, before finetuning SoftNet+FSO on CUB-200-2011, we pre-trained WSN+FSO on the ImageNet-1K dataset, as shown in Table V. The WSN+FSO outperforms WSN and baselines.
Table V: Image Classification Performances on ImageNet-1K.
@@
Table VI: (VIL), PSNR results with Fourier Subnueral Operator (FSO) layer ($f$-NeRVnormal-$\ast$) on UVG17 Video Sessions with average PSNR and Backward Transfer (BWT) of PSNR. Note that $\ast$ denotes our reproduced results.
@@ 7.1 Comparisons with Baselines
Peak signal-to-noise ratio (PSNR). To compare WSN+FSO with conventional representative continual learning methods such as EWC, iCaRL, ESMER, and WSN+FSO, we prepare the reproduced results, as shown in Table VI. The architecture-based WSN outperformed the regularized method and replay method. The sparseness of WSN does not significantly affect sequential video representation results on two sequential benchmark datasets. Our WSN+FSO outperforms all conventional baselines including WSN and MLT (upper-bound of WSN) on the UVG17 benchmark datasets. Moreover, our performances of WSN with $f$-NeRV3 are better than those of $f$-NeRV2 since $f$-NeRV3 tends to represent local textures, stated in the following Section 7.2. Note that the number of parameters of MLT is precisely the same as those of WSN.
@@ 7.4 Ablation Studies of WSN+FSO
Variations of FSO. We prepared several ablation studies to prove the effectiveness of FSO. First, we show the performances of only real part (ignore an imaginary part) in f-NeRV2/3 as shown in Table VII. The PSNR performances of only real part were lower than those of both real and imaginary parts in f-NeRV2/3. We infer that the imaginary part of the winning ticket improves the implicit neural representations. Second, we also investigate the effectiveness of only FSO without Conv. Layer in f-NeRV2/3, as shown in Table VIII. The PSNR performances were lower than FSO with Conv block. Therefore, the ensemble of FSO and Conv improves the implicit representations. Lastly, we investigate the effectiveness of sparse FSO in STL, as shown in Table IX. The sparse FSO boots the PSNR performances in STL. These ablation studies further strengthen the effectiveness of FSO for sequential neural implicit representations.
@@
Table IX: (VIL), WSN+Fourier Subnueral Operator (FSO) layer ($f$-NeRVnormal-$\ast$) on UVG8 Video Sessions with average PSNR and Backward Transfer (BWT).

### v4

@@ 3.3.2 Soft-Subnetworks (SoftNet) with FSO
3.4 Optimization for TIL, TaIL, VIL, and FSCIL
3.4.1 Winning SubNetworks (WSN) for TIL and TaIL
@@
Algorithm 1 Training WSNs for TIL, TaIL, and VIL.
@@ 3.4.3 Soft-SubNetworks (SoftNet) for FSCIL
17:       $\bm{\theta}_{\ast}\leftarrow\bm{\theta}_{\ast}-\beta\left(\frac{\partial\mathcal{L}}{\partial\bm{\theta}_{\ast}}\odot\bm{m}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}minor}}\right)$
@@ 4 Experiments
We validate our method on several benchmark datasets against relevant continual learning baselines on Task-Incremental Learnings (TIL and TaIL, Section 3.4.1), Video Incremental Learning (VIL, Section 3.4.2), and Few-shot Class Incremental Learning (FSCIL, Section 3.4.3).
@@ 4.1 Task-incremental Learning (TIL)
Experimental settings. As we directly implement our method from the official code of [15], we provide the values for HAT and GPM reported in [15]. For Omniglot Rotation and Split CIFAR-100 Superclass, we deploy the proposed architecture in multi-head settings with hyperparameters as reported in [25]. All our experiments run on a single-GPU setup of NVIDIA V100. We evaluate all methods based on the following two metrics:
@@
    Accuracy (ACC) measures the average of the final classification accuracy on all tasks: $\mathrm{ACC}=\frac{1}{\mathcal{T}}\sum_{i=1}^{\mathcal{T}}A_{\mathcal{T},i}$, where $A_{\mathcal{T},i}$ is the test accuracy for task $i$ after training on task $\mathcal{T}$.
@@
    Backward Transfer (BWT) measures the forgetting during continual learning. Negative BWT means that learning new tasks causes the forgetting of past tasks: $\mathrm{BWT}=\frac{1}{\mathcal{T}-1}\sum_{i=1}^{\mathcal{T}-1}A_{\mathcal{T},i}-A_{i,i}$.
@@
4.2 Task-agnositic Incremental Learning (TaIL)
Datasets. We evaluate our WSN+FSO on three popular datasets: Seq-CIFAR10 [37], Seq-CIFAR100 [103], and Seq-TinyImageNet [104]. Seq-CIFAR10 comprises 5 disjoint tasks containing 2 classes and 10k training samples. Seq-CIFAR100 consists of 5 disjoint tasks, each with 20 classes and 10k training samples. Seq-TinyImageNet includes 10 disjoint tasks, each with 20 classes and 10k training samples. Detailed statistics for these datasets can be found in [40]. All experiments address task-agnostic problems where no task ID is provided during training and testing.
Baseline. We compare our FSO with baselines (replay-based methods, Finetune, WSN, and WSN+FSO, Joint) under the experimental setting [40], as shown in Table II. Here, Joint (Upper-bound) denotes the method that all the tasks jointly while Finetune (Lower-bound) denotes the method that learns all tasks sequentially without any memory buffers. Additionally, we compare WSN+FSO with replay-based continual learning methods that maintain a single learning model to perform continual learning (without keeping the extra model [40]). Note that FSO is used at the 3th residual blocks of ResNet18 for TaIL.
Training and Testing. In training time, we follow the continual learning methods [41, 40] with standard ResNet18 for all the task-agnostic experiments. We also use stochastic gradient descent (SGD) to optimize the parameters of WSN+FSO. The batch size is set to 32 for fair comparisons with the prior works. For our experiment results, we report the average and standard deviation of the mean test accuracy of all the sessions across 5 runs with different seeds. At the last epoch of the $s$-th session, we obtain the current subnetwork for the current session and save the subnetwork sequentially so that we have $s$ numbers of subnetworks. Note that we follow the subnetwork’s session identification algorithm in the test for TaIL, as stated in Section 3.4.
@@ 4.3 Few-shot Class Incremental Learning (FSCIL)
Table I: (TIL), Performance comparisons of the proposed method and other baselines - PackNet [23] and SupSup [20] - on various benchmark datasets. We report the mean and standard deviation of the average accuracy (ACC) and average backward transfer (BWT) across $5$ independent runs with five seeds under the same experimental setup [33]. $\dagger$    denotes results reported from [33].
@@
La-MaML [102] 71.37 ($\pm$ 0.7) -5.39 ($\pm$ 0.5) 54.44 ($\pm$ 1.4) -6.65 ($\pm$ 0.9) 66.90 ($\pm$ 1.7) -9.13 ($\pm$ 0.9)
GPM [15] 73.18 ($\pm$ 0.5) -1.17 ($\pm$ 0.3) 57.33 ($\pm$ 0.4) -0.37 ($\pm$ 0.1) 67.39 ($\pm$ 0.5)  1.45 ($\pm$ 0.2)
FS-DGPM [33] 74.33 ($\pm$ 0.3) -2.71 ($\pm$ 0.2) 58.81 ($\pm$ 0.3) -2.97 ($\pm$ 0.4) 70.41 ($\pm$ 1.3) -2.11 ($\pm$ 0.9)
@@
MTL (Upper-bound)   79.75 ($\pm$ 0.4) -   61.00 ($\pm$ 0.2) -   77.10 ($\pm$ 1.1) -
@@ 4.4 Video incremental Learning (VIL)
$\begin{split}\bm{\Gamma}(s,t)=&[\leavevmode\nobreak\ \sin(b^{0}\pi s),\cos(b^{0}\pi s),\cdots,\sin(b^{l-1}\pi s),\cos(b^{l-1}\pi s),\\ &\leavevmode\nobreak\ \leavevmode\nobreak\ \sin(b^{0}\pi t),\cos(b^{0}\pi t),\cdots,\sin(b^{l-1}\pi t),\cos(b^{l-1}\pi t)\leavevmode\nobreak\ ],\end{split}$
where the hyperparameters are set to $b=1.25$ and $l=80$ such that $\bm{\Gamma}(s,t)\in\mathbb{R}^{1\times 160}$. As differences from the previous NeRV model, the first layer of the MLP has its input size expanded from 80 to 160 to incorporate both frame and video indices, and distinct head layers after the NeRV block are utilized for each video. For the loss objective in Equation 8, $\alpha$ is set to $0.7$. We evaluate the video quality, average video session quality, and backward transfer with PSNR.
@@
VIL’s performance metrics We evaluate all methods based on the following continual learning metrics:
@@
    Average Peak signal-to-noise ratio (PSNR) measures the average of the final performances on all video sessions: $\mathrm{PSNR}=\frac{1}{N}\sum_{s=1}^{N}A_{N,s}$, where $A_{N,s}$ is the test PSNR for session $s$ after training on the final video session $S$.
@@ 5.1 Comparisons with baselines in TIL
We use a multi-head setting to evaluate our WSN algorithm under the more challenging visual classification benchmarks. The WSN+FSO’s performances are compared with others in terms of two measurements on three major benchmark datasets as shown in Table I. Our WSN+FSO outperformed all state-of-the-art, achieving the best average accuracy of 79.00%, 61.70%, and 72.04%. WSN+FSO is also a forget-free model (BWT = ZERO), aligned with architecture-based models such as PackNet, SupSup, and WSN in these experiments. In addition, to show the effectiveness of the large single-scale task performance, we prepare WSN+FSO trained on the ImageNet-1K dataset, as shown in Table VI. WSN+FSO outperformed all baselines.
@@ 5.2 Statistics of WSN+FSO’s Representations in TIL
6 Results On Task-agnostic IL (TaIL)
Performances. We set a baseline as WSN to compare with other SOTA methods, in the Task-agnostic Incremental Learning (TaIL) scenario. In seq-CIFAR10 and seq-CIFAR100, WSN+FSO (without any buffer sample) outperformed all baselines and the upper bound (MTL), as shown in Table II. Concretely, FSO’s global abstraction power in the residual block led to the best performances across all agnostic tasks, as shown in Figure 6. In seq-TinyImageNet, the performance of WSN+FSO is lower than that of LODE (DER++). In this setting, Loss decouples learning (LODE) and augmentation methods (DER++) led to close to joint training results. We would focus on how image augmentation and loss decoupling methods affect the selection of subnetworks in continual learning scenarios as future works.
Table II: (TaIL), Incremental Classification results (without keeping the extra model [40]), which are averaged across 5 runs with the different seeds.
      Seq-CIFAR10       Seq-CIFAR100       Seq-TinyImageNet
Buffer Size 500 5120 500 5120 500 5120
SCR [35] 57.95 ±$\pm 1.57$ 82.47 ±$\pm 0.44$ 23.06 ±$\pm 0.22$ 45.02 ±$\pm 0.67$ 8.37 ±$\pm 0.26$ 18.20 ±$\pm 0.48$
PCR [36] 65.74 ±$\pm 3.29$ 82.58 ±$\pm 0.42$ 28.38 ±$\pm 0.46$ 52.51 ±$\pm 1.61$ 11.88 ±$\pm 1.61$ 26.39 ±$\pm 1.64$
MIR [37] 63.93 ±$\pm 0.39$ 83.73 ±$\pm 0.97$ 27.80 ±$\pm 0.52$ 53.73 ±$\pm 0.82$ 11.22 ±$\pm 0.43$ 30.60 ±$\pm 0.40$
ER-ACE [38] 68.45 ±$\pm 1.78$ 83.49 ±$\pm 0.40$ 40.67 ±$\pm 0.06$ 58.56 ±$\pm 0.91$ 17.73 ±$\pm 0.56$ 37.99 ±$\pm 0.17$
ER [39] 61.78 ±$\pm 0.72$ 83.64 ±$\pm 0.95$ 27.69 ±$\pm 0.58$ 53.86 ±$\pm 0.57$ 10.36 ±$\pm 0.11$ 27.54 ±$\pm 0.30$
LODE (ER) [40] 68.87 ±$\pm 0.71$ 83.73 ±$\pm 0.48$ 41.52 ±$\pm 1.22$ 58.59 ±$\pm 0.48$ 17.77 ±$\pm 1.03$ 38.34 ±$\pm 0.04$
DER++ [41] 73.29 ±$\pm 0.96$ 85.66 ±$\pm 0.14$ 42.08 ±$\pm 1.71$ 62.73 ±$\pm 0.58$ 19.28 ±$\pm 0.61$ 39.72 ±$\pm 0.47$
LODE (DER++) [40] 75.45 ±$\pm 0.90$ 85.78 ±$\pm 0.40$ 46.31 ±$\pm 1.01$ 64.00 ±$\pm 0.48$ 21.15 ±$\pm 0.68$ 40.31 ±$\pm 0.03$
Finetune (Lower-bound)           19.65 ±$\pm 0.03$           17.41 ±$\pm 0.09$            8.13 ±$\pm 0.04$
WSN, c=70.0%           94.67 ±$\pm 0.90$           46.24 ±$\pm 0.60$           18.98 ±$\pm 0.17$
WSN, c=70.0% + FSO           94.90 ±$\pm 0.50$           71.12 ±$\pm 0.33$           20.90 ±$\pm 0.21$
MTL (Upper-bound)           91.86 ±$\pm 0.26$           70.10 ±$\pm 0.60$           59.82 ±$\pm 0.31$
Refer to caption Refer to caption
(a) Seq-CIFAR10 (b) Seq-CIFAR100
Figure 6: (TaIL), the accuracy on Seq-CIFAR10 and Seq-CIFAR100.
@@ 7 Results On Few-shot CIL (FSCIL)
Table III: (FSCIL), Classification accuracy of ResNet18 on CIFAR-100 for 5-way 5-shot incremental learning with the same class split as in TOPIC [109]. denotes the results reported from [96].
@@
Table IV: (FSCIL), Classification accuracy of ResNet18 on miniImageNet for 5-way 5-shot incremental learning with the same class split as in TOPIC [109]. denotes results reported from [96].
@@
Table V: (FSCIL), Classification accuracy of ResNet18 on CUB-200-2011 for 10-way 5-shot incremental learning (TOPIC class split [105]). denotes results reported from [96].
@@
Comparisons with SOTA. We compare SoftNet+FSO with the following state-of-art-methods on TOPIC class split [105] of three benchmark datasets - CIFAR100 (Table III), miniImageNet (Table IV), and CUB-200-2011 (Table V). Leveraged by regularized backbone ResNet, SoftNet+FSO outperformed all existing architecture baselines FSLL [91], FACT [116], and WaRP [117] on CIFAR100, miniImageNet. On CUB-200-201, the performances of SoftNet+FSO are comparable with those of ALICE and LIMIT, considering that ALICE used class/data augmentations and LIMIT added an extra multi-head attention layer. Note NC-FSCIL [120] could not be comparable with SoftNet+FSO since it focuses mainly on replaying prototype-based classifiers rather than backbone representations to obtain balanced categorical prototypes. Lastly, before finetuning SoftNet+FSO on CUB-200-2011, we pre-trained WSN+FSO on the ImageNet-1K dataset, as shown in Table VI. The WSN+FSO outperforms WSN and baselines.
Table VI: Image Classification Performances on ImageNet-1K.
@@
[…]
@@ 8.1 Comparisons with Baselines
Video Representations. To compare WSN+FSO with conventional representative continual learning methods such as EWC, iCaRL, ESMER, and WSN+FSO, we prepare the reproduced results, as shown in Table VII. The architecture-based WSN outperformed the regularized method and replay method. The sparseness of WSN does not significantly affect sequential video representation results on two sequential benchmark datasets. Our WSN+FSO outperforms all conventional baselines including WSN and MTL (upper-bound of WSN) on the UVG17 benchmark datasets. Moreover, our performances of WSN with $f$-NeRV3 are better than those of $f$-NeRV2 since $f$-NeRV3 tends to represent local textures, stated in the following Section 8.2. Note that the number of parameters of MLT is precisely the same as those of WSN.
@@ 8.4 Ablation Studies of WSN+FSO
Variations of FSO. We prepared several ablation studies to prove the effectiveness of FSO. First, we show the performances of only real part (ignore an imaginary part) in f-NeRV2/3 as shown in Table VIII. The PSNR performances of only real part were lower than those of both real and imaginary parts in f-NeRV2/3. We infer that the imaginary part of the winning ticket improves the implicit neural representations. Second, we also investigate the effectiveness of only FSO without Conv. Layer in f-NeRV2/3, as shown in Table IX. The PSNR performances were lower than FSO with Conv block. Therefore, the ensemble of FSO and Conv improves the implicit representations. Lastly, we investigate the effectiveness of sparse FSO in STL, as shown in Table X. The sparse FSO boots the PSNR performances in STL. These ablation studies further strengthen the effectiveness of FSO for sequential neural implicit representations.
@@
[…]